# Frieren: Efficient Video-to-Audio Generation with Rectified Flow Matching

### v2

@@ Abstract
Video-to-audio (V2A) generation aims to synthesize content-matching audio from silent video, and it remains challenging to build V2A models with high generation quality, efficiency, and visual-audio temporal synchrony. We propose Frieren, a V2A model based on rectified flow matching. Frieren regresses the conditional transport vector field from noise to spectrogram latent with straight paths and conducts sampling by solving ODE, outperforming autoregressive and score-based models in terms of audio quality. By employing a non-autoregressive vector field estimator based on a feed-forward transformer and channel-level cross-modal feature fusion with strong temporal alignment, our model generates audio that is highly synchronized with the input video. Furthermore, through reflow and one-step distillation with guided vector field, our model can generate decent audio in a few, or even only one sampling step. Experiments indicate that Frieren achieves state-of-the-art performance in both generation quality and temporal alignment on VGGSound, with alignment accuracy reaching 97.22%, and 6.2% improvement in inception score over the strong diffusion-based baseline. Audio samples are available at http://frieren-v2a.github.io.
@@ 1 Introduction
Recent advancements in deep generative models have brought revolutionary enhancements to the quality and diversity of AI-generated content, including text [25], images [27, 28, 1], videos [30, 21] and audios [17, 18]. Among various content-generation tasks, video-to-audio (V2A) generation aims to synthesize semantically relevant and temporally aligned audio from video frames. Due to its immense potential for application in film dubbing, game development, YouTube content creation and other areas, the task of V2A has attracted widespread attention.
@@
We name our model Frieren for efficient video-to-audio generation framework with rectfied flow matching. Experiments indicate that Frieren outperforms strong baselines in terms of audio quality, generation efficiency, and temporal alignment on VGGSound [2], achieving a 6.2% improvement in inception score (IS) and a generation speed 7.3$\times$ that of Diff-Foley, as well as temporal alignment accuracy of up to 97.22% in 25 steps. Additionally, Frieren combining reflow and distillation achieves alignment accuracy of up to 97.85% with just one step, with a 9.3$\times$ acceleration compared to 25-step sampling, further boosting generation efficiency.
@@ 3.2 Model architecture
We conduct experiments with two types of visual representations. For a fair comparison with Diff-Foley [23], we mainly utilize the CAVP feature proposed in [23], which is a visual-audio contrastive feature considering both content and temporal alignment. Meanwhile, to investigate the impact of visual feature characteristics on model performance, we also attempt the visual feature from MAViL [12], which is an advanced self-supervised visual-audio representation learner that employs both masked-reconstruction and contrastive learning, and exhibits formidable performance in audio-visual understanding (See section 4.3.2 for comparison). For audio representation, we follow a previous text-to-audio work [11] to train a mel-spectrogram VAE with 1D convolution over the temporal dimension. Details of the VAE are provided in appendix A
@@ 4 Experiments
DDPM 11.79 10.09 2.86 1.77 2.36 95.33 3.72 $\pm$ 0.05 3.79 $\pm$ 0.07
Frieren (Ours) 12.25 12.42 2.73 1.32 2.49 97.22 3.78 $\pm$ 0.06 3.90 $\pm$ 0.05
@@ 4.1 Experiment setup
We combine objective and subjective metrics to evaluate model performance over audio quality, diversity, and temporal alignment. For objective evaluation, we calculate Frechet distance (FD), inception score (IS), Kullback–Leibler divergence (KL), Frechet audio distance (FAD), kernel inception distance (KID), and alignment accuracy (Acc). We utilize audio evaluation tools provided by AudioLDM [17], which are widely used in audio generation tasks, as well as the alignment classifier provided in [23]. For subjective evaluation, we conduct crowd-sourced human evaluations with 1-5 Likert scales and report mean-opinion-scores (MOS) over audio quality (MOS-Q) and content alignment (MOS-A) with 95% confidence intervals (CI). We sample 10 audios for each test video for evaluation. Details of subjective evaluation are provided in appendix B.
@@
We adopt three advanced V2A models as baselines, including: 1) SpecVQGAN [13], a transformer-based autoregressive model generating spectrogram VQVAE indices from visual features; 2) Im2Wav [29], a hierarchical autoregressive V2A model predicting audio VQVAE indices conditioned on CLIP features; and 3) Diff-Foley [23], a strong latent-diffusion-based V2A model. For SpecVQGAN, we evaluate two versions using RGB+Flow and ResNet features as input visual conditions. For Diff-Foley, we evaluate its performance with and without classifier guidance to examine the impact of its complex external alignment mechanism. To better validate the superiority of our rectified flow model, we also train a diffusion model sharing the same architecture as Frieren but has a different prediction target, labeled as DDPM in the following tables. We use DPM-Solver [22] for all diffusion models, and Euler method 5 for our rectified flow model.
@@ 4.2 Results and analysis
The results of different models are illustrated in table 1. We sample with diffusion models and Frieren with 25 steps, and report the result of Frieren without reflow and distillation, which shows the best overall performance with a high number of sampling steps. It can be seen that Frieren significantly outperforms other models in IS, FAD, and alignment accuracy, with the values reaching up to 12.42, 1.32, and 97.22%, together with the highest subjective scores of 3.78 and 3.90 on quality and alignment. For FD, KL, and KID, the scores of Frieren are also very close to the best values, with differences of only 0.46, 0.16, and $\times 10^{-3}$. This indicates the effectiveness of our approach. Generally, the performance of Frieren surpasses that of the DDPM, demonstrating the superiority of rectified flow matching over the score-based paradigm of diffusion. Additionally, both Frieren and DDPM outperform Diff-Foley in temporal alignment, proving that our architecture design achieves strong temporal alignment without the need for complex mechanisms, and can produce audio that is highly synchronized with visual input. Additionally, our model also has an advantage in sampling time, with details provided in appendix C.
Table 2: Results of Frieren and the diffusion baselines under different sampling steps. CG denotes classifier guidance, R denotes reflow and D denotes one-step distillation.
@@
DDPM 11.79 10.09 2.86 1.77 2.36 95.33 3.72 $\pm$ 0.05 3.79 $\pm$ 0.07
Frieren (R ✗, D ✗) 12.25 12.42 2.73 1.32 2.49 97.22 3.78 $\pm$ 0.06 3.90 $\pm$ 0.05
Frieren (R ✓, D ✗) 13.38 9.79 2.64 1.65 2.70 97.36 3.61 $\pm$ 0.07 3.88 $\pm$ 0.05
Figure 3: One-step generation results of different models. (a): The content of the input video is a woman playing the violin. (b): Diff-Foley generates meaningless audio with one step. (c, d): DDPM and Frieren without reflow generate highly noisy audio. (e, f): reflow enables Frieren to generate meaningful audio in one step, and distillation further improves the one-step generation quality.
@@
We further demonstrate the results of diffusion models and Frieren on reduced sampling steps in table 2 to illustrate the impact of reflow and one-step distillation, together with trend graphs of IS and FAD in figure 4 for intuitive presentation. The data for reflow are generated with the Euler method for 25 steps. We observe an obvious drop in performance of the diffusion baselines as well as Frieren without reflow when sampling with as few as 5 steps, and their scores become extremely poor when we further reduce the step number to 1. Figure 3 (b) (c) and (d) illustrate that the audio generated by these models degrades into unacceptably noisy or meaningless audio within one step. This is due to the convoluted nature of the sampling trajectories of these models, which disables them from sampling with large step sizes and few steps. We also notice that when sampling with 5 steps, using additional classifier guidance deteriorates the audio quality and synchrony of Diff-Foley, where alignment accuracy and IS drop by 18.0% and 1.72 respectively, while FD, KL, and KID increase by 9.47, 0.17, and 0.94 ×$\times 10^{-3}$. This indicates the lack of robustness of the complex alignment mechanism that Diff-Foley relies on.
In contrast, Frieren with reflow achieves an alignment accuracy of up to 96.82% in just 5 steps, with significant advantages in quality, diversity, and subjective metrics. Additionally, it maintains an accuracy of 94.96% in single-step generation, as well as decent quality and diversity. This proves that reflow functions significantly in straightening the sampling trajectories, enabling the rectified flow model to generate decent audio with a small number of sampling steps. Furthermore, single-step distillation following reflow further improves the model performance with one step, with alignment accuracy reaching up to 97.85%, and KL, FAD, and KID being very close to the 25-step results of Frieren trained once, with differences of only 0.08, 0.2 and 2.1×$\times 10^{-3}$. It also achieves high MOS-Q and MOS-A of 3.48 and 3.93.
@@ 4.3.1 Model size of vector field estimator
We observe that when the model parameters are reduced to 71M, performance declines across all metrics, where FD, KL, FAD, and KID increase by 0.77, 0.05, 0.18, and $\times 10^{-3}$, and IS, alignment accuracy, MOS-Q and MOS-A drop by 0.26, 1.18%, 0.07 and 0.07, respectively. However, when the parameter number increases to 421M, there is a performance degradation across multiple metrics, with KL, FAD, and KID increasing by 0.03, 0.04, and 0.48×$\times 10^{-3}$, and IS, alignment acc declining by 0.13 and 2.06%. We speculate that this anomalous phenomenon may be due to the convergence difficulty for the larger model under similar training steps, or the redundant model capacity tends to cause overfitting on a relatively small dataset like VGGSound, deteriorating the model’s generalization performance. In summary, we achieve relatively balanced model performance with the parameter of the estimator being around 160M. Details of model parameters are provided in appendix A.
@@ 4.3.2 Visual feature characteristics
In table 4, we compare the results of Frieren using two different types of visual features from CAVP and MAViL. Intuitively, the MAViL feature should be more robust and contain richer audio-related semantic information, as it utilizes masked-reconstruction together with inter-modal and intra-modal contrastive learning, in contrast to CAVP trained solely with inter-modal contrastive learning. On the other hand, however, due to MAViL’s convolutional downsampling in the temporal dimension, its feature sequence has a lower effective FPS of 2 with the same 4 FPS video input as CAVP. The results in the table indicate that the model with MAViL feature excels in audio diversity, with differences of FD, KL, and FAD being 0.17, 0.24, and 0.06. Meanwhile, it exhibits a 7.05% decrease in alignment accuracy and a 0.25 decrease in IS. This result yields two insights for V2A tasks: 1) at relatively low frame rates, the frame rate of features, rather than content, is more likely to become the bottleneck for audio quality and visual-audio synchrony; 2) compared to high video frame rates, the semantic information and robustness of visual features are more crucial for the diversity of generated audio.
@@ 4.3.4 Re-weighting RFM objective
We conduct ablation on RFM objective re-weighting and report the results in table 5. We can see that compared to the vanilla objective, introducing re-weighting results in improvements of 0.22 and 0.18% for IS and alignment accuracy. This validates the positive impact of objective re-weighting on audio quality and temporal alignment. On the other hand, objective re-weighting causes a decrease in audio diversity, with differences in FD, FAD, and KID being 0.30, 0.07, and 0.37×$\times 10^{-3}$, respectively.
@@ 5 Conclusion
Appendix A Model parameters and architecture details
@@ Appendix B Subjective evaluation
For each evaluated model, we randomly select 150 items for subjective evaluation, accounting for about 1% of the entire test split.
@@ Appendix C Time efficiency
Frieren (step=25) 0.288
Frieren (step=5) 0.064
Frieren (step=1) 0.031
In table 8, we compare the inference time per sample of different models. The inference is conducted on a single RTX-4090 GPU with a batch size of 1. We can see that the inference procedure of transformer-based autoregressive models, including SpecVQGAN and Im2Wav, is more time-consuming, especially for Im2Wav, which takes several minutes to generate a single sample. This is because Im2Wav conducts a cascaded generation with 2 transformers. Moreover, its use of high-bitrate audio VQVAE results in very long sequences of audio representation, significantly increasing the inference time required for the transformers, which has quadratic time complexity concerning sequence length. In contrast, Diff-Foley and Frieren require less inference time, and Frieren enjoys a higher speed, achieving 7.3 times faster than Diff-Foley with 25 sampling steps. This is the result of a combination of multiple factors, including model architecture, model parameters, sampling methods, and so on. Furthermore, when using Frieren model with reflow and one-step distillation, we can generate 5-step sampled audio in 0.064 seconds and 1-step sampled audio in just 0.031 seconds, achieving 4.5$\times$ and 9.3$\times$ acceleration compared to 25-step sampling. This demonstrates the extremely high generation efficiency of our model on the task of V2A.
Appendix D Limitations and boarder impacts

### v3

# Frieren: Efficient Video-to-Audio Generation Network with Rectified Flow Matching

@@ Abstract
Video-to-audio (V2A) generation aims to synthesize content-matching audio from silent video, and it remains challenging to build V2A models with high generation quality, efficiency, and visual-audio temporal synchrony. We propose Frieren, a V2A model based on rectified flow matching. Frieren regresses the conditional transport vector field from noise to spectrogram latent with straight paths and conducts sampling by solving ODE, outperforming autoregressive and score-based models in terms of audio quality. By employing a non-autoregressive vector field estimator based on a feed-forward transformer and channel-level cross-modal feature fusion with strong temporal alignment, our model generates audio that is highly synchronized with the input video. Furthermore, through reflow and one-step distillation with guided vector field, our model can generate decent audio in a few, or even only one sampling step. Experiments indicate that Frieren achieves state-of-the-art performance in both generation quality and temporal alignment on VGGSound, with alignment accuracy reaching 97.22%, and 6.2% improvement in inception score over the strong diffusion-based baseline. Audio samples and code are available at http://frieren-v2a.github.io.
@@ 1 Introduction
Recent advancements in deep generative models have significantly enhanced the quality and diversity of AI-generated content, including text [27], images [29, 30, 1], videos [32, 23] and audios [19, 20]. Among various content-generation tasks, video-to-audio (V2A) generation aims to synthesize semantically relevant and temporally aligned audio from video frames. Due to its immense potential for application in film dubbing, game development, YouTube content creation and other areas, the task of V2A has attracted widespread attention.
@@
We name our model Frieren for efficient video-to-audio generation network with rectfied flow matching. Experiments indicate that Frieren outperforms strong baselines in terms of audio quality, generation efficiency, and temporal alignment on VGGSound [2], achieving a 6.2% improvement in inception score (IS) and a generation speed 7.3$\times$ that of Diff-Foley, as well as temporal alignment accuracy of up to 97.22% in 25 steps. Additionally, Frieren combining reflow and distillation achieves alignment accuracy of up to 97.85% with just one step, with a 9.3$\times$ acceleration compared to 25-step sampling, further boosting generation efficiency.
@@ 3.2 Model architecture
Various audio-aligned visual representations [9, 25, 14, 38, 37, 36, 39] can potentially be applied to video-to-audio generation, and we conduct experiments with two types of visual representations. For a fair comparison with Diff-Foley [25], we mainly utilize the CAVP feature proposed in [25], which is a visual-audio contrastive feature considering both content and temporal alignment. Meanwhile, to investigate the impact of visual feature characteristics on model performance, we also attempt the visual feature from MAViL  [^1] Implementation of MAViL is from av-superb [33]: https://github.com/roger-tseng/av-superb [14], which is an advanced self-supervised visual-audio representation learner that employs both masked-reconstruction and contrastive learning, and exhibits formidable performance in audio-visual understanding (See section 4.3.2 for comparison). For audio representation, we follow a previous text-to-audio work [13] to train a mel-spectrogram VAE with 1D convolution over the temporal dimension. Details of the VAE are provided in appendix A.
@@ 4 Experiments
LDM 11.79 10.09 2.86 1.77 2.36 95.33 3.72 $\pm$ 0.05 3.79 $\pm$ 0.07
Frieren 12.26 12.42 2.73 1.32 2.49 97.22 3.78 $\pm$ 0.06 3.90 $\pm$ 0.05
Frieren (Dopri5) 11.64 12.76 2.75 1.37 2.39 96.87 3.81 $\pm$ 0.06 3.85 $\pm$ 0.06
@@ 4.1 Experiment setup
We combine objective and subjective metrics to evaluate model performance over audio quality, diversity, and temporal alignment. For objective evaluation, we calculate Frechet distance (FD), inception score (IS), Kullback–Leibler divergence (KL), Frechet audio distance (FAD), kernel inception distance (KID), and alignment accuracy (Acc). We utilize audio evaluation tools provided by AudioLDM [19], which are widely used in audio generation tasks, as well as the alignment classifier provided in [25]. For metrics with reference like FAD, we duplicate the reference audio samples in the test set for 10 times as we generate 10 samples for each data item. For subjective evaluation, we conduct crowd-sourced human evaluations with 1-5 Likert scales and report mean-opinion-scores (MOS) over audio quality (MOS-Q) and content alignment (MOS-A) with 95% confidence intervals (CI). We sample 10 audios for each test video for evaluation. Details of subjective evaluation are provided in appendix B.
@@
We adopt three advanced V2A models as baselines, including: 1) SpecVQGAN [15], a transformer-based autoregressive model generating spectrogram VQVAE indices from visual features; 2) Im2Wav [31], a hierarchical autoregressive V2A model predicting audio VQVAE indices conditioned on CLIP features; and 3) Diff-Foley [25], a strong latent-diffusion-based V2A model. For SpecVQGAN, we evaluate two versions using RGB+Flow and ResNet features as input visual conditions. For Diff-Foley, we evaluate its performance with and without classifier guidance to examine the impact of its complex external alignment mechanism. To better validate the superiority of our rectified flow model, we also train a diffusion model sharing the same architecture as Frieren but has a different prediction target, labeled as LDM in the following tables. For diffusion models, we use DPM-Solver [24] for sampling. For our rectified flow model, we use the Euler method 5 in most cases without further specification. We also explore the more advanced Dormand–Prince method (Dopri5) [4] method for higher generation quality.
@@ 4.2 Results and analysis
The results of different models are illustrated in table 1. We sample with diffusion models and Frieren with 25 steps, and report the result of Frieren without reflow and distillation, which shows the best overall performance with a high number of sampling steps. It can be seen that Frieren significantly outperforms other models in IS, FAD, and alignment accuracy, with the values reaching up to 12.42, 1.32, and 97.22%, together with high subjective scores of 3.78 and 3.90 on quality and alignment. For FD, KL, and KID, the scores of Frieren are also very close to the best values among other models. When we employ the higher-order Dopri5 ODE solver, Frieren achieves further improvements in FD and IS, attaining best values of 11.64 and 12.76, respectively, while maintaining stable performance in other objective metrics, at the cost of slower sampling speed. This indicates the effectiveness of our approach. Generally, the performance of Frieren surpasses that of the LDM, demonstrating the superiority of rectified flow matching over the score-based paradigm of diffusion. Additionally, both Frieren and LDM outperform Diff-Foley in temporal alignment, proving that our architecture design achieves strong temporal alignment without the need for complex mechanisms, and can produce audio that is highly synchronized with visual input. Additionally, our model also has an advantage in sampling time, with details provided in appendix C.
Table 2: Results of Frieren and Diff-Foley under different sampling steps. CG denotes classifier guidance, R denotes reflow and D denotes one-step distillation.
@@
Frieren (R ✗, D ✗) 12.26 12.42 2.73 1.32 2.49 97.22 3.78 $\pm$ 0.06 3.90 $\pm$ 0.05
Frieren (R ✓, D ✗) 13.39 9.79 2.64 1.66 2.01 97.36 3.61 $\pm$ 0.07 3.88 $\pm$ 0.05
Figure 3: One-step generation results of different models. (a): The content of the input video is a woman playing the violin. (b): Diff-Foley generates meaningless audio with one step. (c, d): LDM and Frieren without reflow generate highly noisy audio. (e, f): reflow enables Frieren to generate meaningful audio in one step, and distillation further improves the one-step generation quality.
@@
We further demonstrate the results of Diff-Foley and Frieren on reduced sampling steps in table 2 to illustrate the impact of reflow and one-step distillation, together with trend graphs of IS and FAD in figure 4 for intuitive presentation. The data for reflow are generated with the Euler method for 25 steps. We observe an obvious drop in performance of Diff-Foley as well as Frieren without reflow when sampling with as few as 5 steps, and their scores become extremely poor when we further reduce the step number to 1. Figure 3 (b) (c) and (d) illustrate that the audio generated by these models as well as LDM degrades into unacceptably noisy or meaningless audio within one step. This is due to the convoluted nature of the sampling trajectories of these models, which disables them from sampling with large step sizes and few steps. We also notice that when sampling with 5 steps, using additional classifier guidance deteriorates the audio quality and synchrony of Diff-Foley, where alignment accuracy and IS drop by 18.0% and 1.72 respectively, while FD, KL, and KID increase by 9.47, 0.17, and 0.94 ×$\times 10^{-3}$. This indicates the lack of robustness of the complex alignment mechanism that Diff-Foley relies on.
@@
In contrast, Frieren with reflow achieves an alignment accuracy of up to 96.82% in just 5 steps, with significant advantages in quality, diversity, and subjective metrics. Additionally, it maintains an accuracy of 94.96% in single-step generation, as well as decent quality and diversity. This proves that reflow functions significantly in straightening the sampling trajectories, enabling the rectified flow model to generate decent audio with a small number of sampling steps. Furthermore, single-step distillation following reflow further improves the model performance with one step, with alignment accuracy reaching up to 97.85%, and KL, FAD, and KID being close to the 25-step results of Frieren trained once, with differences of 0.17, 0.53 and 0.42×$\times 10^{-3}$. It also achieves high MOS-Q and MOS-A of 3.48 and 3.93.
@@ 4.3.1 Model size of vector field estimator
We observe that when the model parameters are reduced to 71M, performance declines across all metrics, where FD, KL, FAD, and KID increase by 0.76, 0.05, 0.18, and $\times 10^{-3}$, and IS, alignment accuracy, MOS-Q and MOS-A drop by 0.26, 1.18%, 0.07 and 0.07, respectively. However, when the parameter number increases to 421M, there is a performance degradation across multiple metrics, with KL, FAD, and KID increasing by 0.03, 0.04, and 0.48×$\times 10^{-3}$, and IS, alignment acc declining by 0.13 and 2.06%. We speculate that this anomalous phenomenon may be due to the convergence difficulty for the larger model under similar training steps, or the redundant model capacity tends to cause overfitting on a relatively small dataset like VGGSound, deteriorating the model’s generalization performance. In summary, we achieve relatively balanced model performance with the parameter of the estimator being around 160M. Details of model parameters are provided in appendix A.
@@ 4.3.2 Visual feature characteristics
In table 4, we compare the results of Frieren using two different types of visual features from CAVP and MAViL. Intuitively, the MAViL feature should be more robust and contain richer audio-related semantic information, as it utilizes masked-reconstruction together with inter-modal and intra-modal contrastive learning, in contrast to CAVP trained solely with inter-modal contrastive learning. On the other hand, however, due to MAViL’s convolutional downsampling in the temporal dimension, its feature sequence has a lower effective FPS of 2 with the same 4 FPS video input as CAVP. The results in the table indicate that the model with MAViL feature excels in audio diversity, with differences of FD, KL, and FAD being 0.18, 0.24, and 0.06. Meanwhile, it exhibits a 7.05% decrease in alignment accuracy and a 0.25 decrease in IS. This result yields two insights for V2A tasks: 1) at relatively low frame rates, the frame rate of features, rather than content, is more likely to become the bottleneck for audio quality and visual-audio synchrony; 2) compared to high video frame rates, the semantic information and robustness of visual features are more crucial for the diversity of generated audio.
@@ 4.3.4 Re-weighting RFM objective
We conduct ablation on RFM objective re-weighting and report the results in table 5. We can see that compared to the vanilla objective, introducing re-weighting results in improvements of 0.22 and 0.18% for IS and alignment accuracy. This validates the positive impact of objective re-weighting on audio quality and temporal alignment. On the other hand, objective re-weighting causes a decrease in audio diversity, with differences in FD, FAD, and KID being 0.31, 0.07, and 0.37×$\times 10^{-3}$, respectively.
@@ 5 Conclusion
Acknowledgment
This work is supported by the National Natural Science Foundation of China under Grant No. 62222211 and No.62072397.
Appendix A Implementation details
@@ Appendix B Subjective evaluation
For each evaluated model, we select 150 items for subjective evaluation, accounting for about 1% of the entire test split.
@@ Appendix C Time efficiency
Frieren (Dopri5, step=25) 1.510
Frieren (Euler, step=25) 0.288
Frieren (Euler, step=5) 0.064
Frieren (Euler, step=1) 0.031
In table 8, we compare the inference time per sample of different models. The inference is conducted on a single RTX-4090 GPU with a batch size of 1. We can see that the inference procedure of transformer-based autoregressive models, including SpecVQGAN and Im2Wav, is more time-consuming, especially for Im2Wav, which takes several minutes to generate a single sample. This is because Im2Wav conducts a cascaded generation with 2 transformers. Moreover, its use of high-bitrate audio VQVAE results in very long sequences of audio representation, significantly increasing the inference time required for the transformers, which has quadratic time complexity concerning sequence length. In contrast, Diff-Foley and Frieren require less inference time, and Frieren with Euler solver enjoys a higher speed, achieving 7.3 times faster than Diff-Foley with 25 sampling steps. This is the result of a combination of multiple factors, including model architecture, model parameters, sampling methods, and so on. Furthermore, when using Frieren model with reflow and one-step distillation, we can generate 5-step sampled audio in 0.064 seconds and 1-step sampled audio in just 0.031 seconds, achieving 4.5$\times$ and 9.3$\times$ acceleration compared to 25-step sampling. This demonstrates the extremely high generation efficiency of our model on the task of V2A.
Appendix D Impact of the vocoder on model performance
Table 9: Comparison of the performance of Diff-Foley and Frieren using the same vocoder.
[…]
BigVGAN
Diff-Foley (CG ✓) 18.02 10.89 2.88 6.32 5.32
Frieren 12.26 12.42 2.73 1.32 2.49
Griffin-Lim
Diff-Foley (CG ✓) 23.94 11.11 3.38 4.72 9.58
Frieren 28.29 10.67 3.17 3.70 12.30
Different selections of vocoders can significantly impact the performance of various audio generation models. Diff-Foley uses the simple Griffin-Lim method [10] to map spectrograms to waveforms, while Frieren employs the more efficient BigVGAN. To compare the performance of the spectrogram generation models while minimizing the influence of the vocoder, we apply BigVGAN and Griffin-Lim separately to each model. The output from Diff-Foley is converted into an 80-bin mel-spectrogram and then fed into BigVGAN. The number of Griffin-Lim iterations for Frieren is the same as Diff-Foley. The results are shown in table 9.
It can be seen that using BigVGAN for Diff-Foley improves its FD, KL, and KID, indicating its effectiveness. On this basis, Frieren outperforms Diff-Foley across all metrics, with a greater difference than when using Griffin-Lim for both. This further demonstrates that our model is superior to Diff-Foley.
On the other hand, when using Griffin-Lim for both models, despite the performance drop, Frieren still surpasses Diff-Foley in KL and FAD, with FAD showing a significant advantage while maintaining competitive FD and IS values. We speculate that the Griffin-Lim algorithm is so weak that it forms a performance bottleneck, narrowing the performance gap between Frieren and Diff-Foley. Additionally, differences in spectrogram hyperparameters may also lead to a performance gap. Diff-Foley uses 128 frequency bins, more than the 80 bins used by Frieren, allowing it to carry finer-grained information and may give Diff-Foley an advantage when using Griffin-Lim.
Appendix E Limitations and boarder impacts